OpenStudy (anonymous):

-3<2x-5 is less than or equal to 11??

OpenStudy (anonymous):

i'll help you

OpenStudy (anonymous):

@abswag13 Thank you

OpenStudy (anonymous):

what do you want me to do with it?

OpenStudy (anonymous):

theres different things i could do with it

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

solve for the inequality

OpenStudy (anonymous):

-3<2x-5 <=11 This implies -3+5<2x<=11+5 this implies 2<2x<=16 this implies 1<x<=8 so the solution is {x belongs to R such that 1<x<=8} it can also be written as (1,8] @xo-beauty

OpenStudy (anonymous):

@Princer_Jones Thank you!